OpenStudy (anonymous):

How do I factor: 2x^4+x^3-8x^2-x+6?

OpenStudy (anonymous):

do you know remainder and factor theorem?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

I can solve this question through it but need to know whether you know it or not?

OpenStudy (anonymous):

i guess not since I haven't heard of factor theorem.. I mean I know the special ones like difference of perfect square, cubes

OpenStudy (anonymous):

i need steps-- I just don't get it

OpenStudy (anonymous):

2x^4+x^3-8x^2-x+6=0 2x^4+x^3-8x^2-x=-6 x(2x^3+x^2-8x-1)=-6 i don't know if it is correct

OpenStudy (anonymous):

\[f(x)=2x ^{4}+x ^{3}-8x ^{2}-x+6\] Let x be 1, f(1)=0 hence (x-1) is the factor. Do you know this method?

OpenStudy (anonymous):

i've seen it before

OpenStudy (anonymous):

great.As we know that (x-1) is the factor,the equation can be wrote like this: (x-1)(Ax3+Bx2+Cx+D)=2x4+x3−8x2−x+6 please find the value of A,B and C and tell me the answer?

OpenStudy (anonymous):

why is x-1 the factor?

OpenStudy (anonymous):

because when x=1 then f(x)=0 and this proves that (x-1) is the factor.